# Time-dependent Hamiltonians and Geometry of Operators Generated by Them

### v1

@@ I Introduction
The quantity computational complexity of a given quantum state roughly measures the minimum number of basic gates required to build the target state starting from an initial reference state. Among various measures to quantify computational complexity, one of the most commonly used is the one proposed by Nielsen and collaborators Nielsen1 ; Nielsen2 ; Nielsen3 . In this approach, the reference and target states correspond to two points on a curved manifold, and all the possible unitary transformations that connect these two points are assigned with a cost function. According to this geometric formulation, the path that minimises the cost function is nothing but the geodesic on that manifold of the unitary operator, and the length of the geodesic is called the Nielsen complexity (NC). This particular measure of complexity has been explored in significant detail in recent literature not only because it can shed light on fundamental aspects of the information contained in a quantum state Myers1 ; BSS ; KKS ; Myers2 ; HM ; BMS ; BS ; Caceres:2019pgf , but also due to the fact that NC can also be used an important probe of quantum quenches ABHKM ; Gautam:2022gci ; Pal:2022rqq ; Chandran:2022vrw ; Camargo:2018eof , quantum phase transitions LWCLTYGG ; JGS ; XYZ ; SK ; JGS2 ; PPS ; WH ; Roca-Jerat:2023mjs ; Jaiswal:2024yts ; Jayarama:2022xta , and quantum chaos BDKP1 ; BDKP2 ; HJU ; Qu:2022zwq , and has various other applications as well Pal:2021jtm ; Chowdhury:2023iwg ; Adhikari:2021ckk ; Basteiro:2021ene ; Camilo:2020gdf ; Erdmenger:2020sup . Specifically, typical behaviour of circuit complexity has been used in ABHKM ; Camargo:2018eof ; Pal:2022rqq ; Gautam:2022gci ; Chandran:2022vrw to understand the unitary dynamics of an isolated quantum system, when the system is taken out of equilibrium by means of a single or multiple sudden or smooth quenches thereby changing the parameters of the Hamiltonian according to different relevant time scales. It has also been used in various works as an important quantum information theoretic tool to detect the zero temperature quantum phase transition in various contexts and models, where typically the discontinuity of the complexity marks the quantum critical point LWCLTYGG ; JGS ; XYZ ; SK ; JGS2 ; PPS ; WH ; Roca-Jerat:2023mjs ; Jaiswal:2024yts ; Jayarama:2022xta . Apart from the NC, there are also other important measures of circuit complexity of quantum states which differ from Nielsen’s approach in many aspects (e.g., the one introduced in CHMP by using the Fubini-Study metric associated with a pure quantum state), however, in this paper, we shall mostly consider Nielsen’s formulation. For a recent review of different approaches to circuit complexity and its application in holography, we refer to Chapman:2021jbh .
@@
It is important to note that, apart from the given target and reference states, the NC crucially depends on the choice of the cost function associated with each of the unitary paths on the manifold. Though the qualitative nature of NC in a quantum system strongly depends on the choice cost function, in Nielsen’s original formulation, this choice remains arbitrary (for an exposition on different cost functions used in this context, we refer to BMS ). Another built-in aspect of NC is the fact that the unitary complexity geometry associated with NC is either right or left invariant under a global transformation of the unitary operator which determines the geometry. As was pointed out in YK ; YK2 ; YANZK ; YANZK2 , for a physical quantum system, this fact is somewhat unsatisfactory since, e.g., the Nielsen complexity geometry corresponding to two Hamiltonians related by a unitary transformation would be different, and hence unacceptable from a physical point of view since they describe the same quantum system. Based on the assumption that NC and hence the cost function should be invariant under both left and right unitary transformation of the generator associated with a given path on the unitary manifold, the notion of bi-invariant complexity was proposed in YK ; YANZK ; YANZK2 . In this paper we mainly consider this bi-invariant formulation of the NC, and apply it to the Hamiltonians of some well known quantum mechanical systems.
@@
Usually, in the problem of calculating NC (or other notions of circuit complexity), the following situation is considered - the reference and target quantum states are given (along with the chosen cost functional and penalty factors for any gate directions, if any), and one has to minimise the costs associated with all the relevant unitary curves to find the NC. However, in this problem, the instantaneous generator of the unitary curve is not commonly given specific importance because we are only interested in the length of the geodesic curve (i.e., complexity). Consider now a different posing of the above-mentioned problem, which is equally important: the time evolution of an initial reference under the Hamiltonian of a given quantum system. In that case, the unitary operator is the time evolution operator itself, and the target state is actually some time-evolved state for a particular value of time. Of course, any arbitrary target state may not be on the unitary curve generated by the system Hamiltonian, and consequently, those states are not achievable under this protocol. To explain further, suppose, one can choose only two from the three quantities: a reference state, a target state, and Hamiltonian that generates the evolution, and third one is naturally constrained by a unitary evolution. In the standard way of defining NC, the first two (i.e., reference and target states) are given, whereas here we are interested in the scenario where the first and third options (i.e., reference state and system Hamiltonian) are pre-assigned. As a natural consequence of the preceding statement, the nature of the parameter associated with the unitary curve connecting the reference and target state is different from the usual treatment of NC. Specifically, in this paper, we construct a bi-invariant geometry associated with the operators generated by the Hamiltonian of quantum mechanical systems. One of the most important such operators is the time evolution operator itself, and in that case, the parameter along the unitary path is time. In summary, our goal here is to calculate the cost of operators along the time evolution curve of a quantum mechanical system by using the framework of bi-invariant complexity geometry (CG).
@@
In Section V, we generalise the results of section IV for the case of the Hamiltonian of an oscillator whose both mass and frequency changes with time. For this Hamiltonian also, we are able to provide an analytical formula for the regularised cost function by using the Lewis - Riesenfeld theory of time-dependent invariant operators. The fact that such a Hamiltonian is actually an element of the $su(1,1)$ Lie algebra provides a way of directly obtaining the exact form of the time evolution operator using techniques of Lie group theory. Using this exact expression for the time-evolution operator, we show that suitably redefining the time parameter, it is possible to map the time evolution operator of an oscillator with time varying mass and frequency to an oscillator with unit mass but time varying frequency. Furthermore, we show that under the same reparametrisation, the total bi-invariant cost evaluated along the time-evolution curves of these two oscillators can be mapped to each other.
We discuss the significance of our results in section VI, and indicate a few possible future directions. This paper also contains two appendices: in Appendix. A we compute the bi-invariant cost and compare it with other measures, such as the change in Shannon entropy for one of the most common physical situations where time-dependent Hamiltonians are used, namely that of a smooth quantum quench, whereas in Appendix. B we provide some details of the computation of the time-evolution operator associated with Hamiltonians which are elements of the $su(1,1)$ Lie algebra.
@@ II Cost functional and the circuit complexity in Nielsen’s approach
Now we come back to the problem of defining such a geometry uniquely. Besides the above mentioned problem of nonuniqueness of the CG, there is one additional feature of such constructions - the resulting geometry is only right (or left) invariant under a global transformation of the unitary operator. As discussed in detail in YANZK ; YANZK2 ; YK , this property is somewhat problematic for defining complexity associated with a quantum mechanical Hamiltonian, since a unitary transformed system is identical to that of the original system, in the sense that physical observables computed from both of them make identical predictions for the original system and the transformed one. Hence, it is very much desirable that for an arbitrary unitary operator $U(s)$, defined on the CG, the cost function or the norm should be such that it is invariant under the unitary transformation of the Hamiltonian, i.e.,
@@ III Complexity geometry associated with time-independent Hamiltonians
We now discuss the construction of the bi-invariant CG of operators generated by the Hamiltonian of some well-known quantum systems. In this section we consider the case when the Hamiltonian is time-independent, while in the next section, and the rest of the paper, we deal with systems whose Hamiltonians are explicit functions of time. The instantaneous Hamiltonian appearing in Eq. (2) is the ordered Hermitian Hamiltonian operator (which we denote as $\mathbf{H}$ in the followings) of a given quantum system, and this equation is nothing but the Schrodinger equation for the time evolution of the state $\big{|}\psi(t)\big{>}$. Furthermore, the unitary operator of our concern is the time evolution operator generated by the system Hamiltonian. At this point, we stress that we want to evaluate the computational cost along the time evolution curve $\mathcal{U}(t)$, not the NC of a time evolved state (or an operator on the time evolution curve) - this fact makes our analysis different from other works where time evolution of NC corresponding to different quantum systems have been reported BDKP1 ; YK2 ; HJU . Specifically, in our set-up, the parameter $s$ characterising the curves in the space of unitary operators is time $t$.  [^1] In this sense, the motivation of the present paper is similar to the recent work Craps:2022ese , which has obtained an upper bound on the NC by restricting the minimisation procedure described in the previous section within a prescribed family of curves connecting the identity and the target operator.
@@
This cost represents the length on a Riemannian metric, which is nothing but the Cartan-Killing metric corresponding to the Lie group associated with the quantum system RG ; FSS . As we shall see, for the cases we consider in this paper, the Hamiltonian is an element of some Lie algebra, and the above mentioned Lie group is associated with this algebra. Clearly, to find out the CG using this cost, we need to calculate the trace of the squared Hamiltonian. However, the Hilbert spaces associated with the systems we are going to consider are infinite dimensional, and hence the trace can be divergent. A procedure of obtaining the CG by suitably regularizing the trace of the Hamiltonian operator using the Hurwitz $\zeta$ functions was described in YK . For later purposes, we first briefly review this procedure and its application to find out the CG of a generalised oscillator, as worked out in YK . We can also provide an alternative interpretation of this regularization in terms of a constant driving force present in the original Hamiltonian (see subsection III.2).
@@
To illustrate the above procedure, the quantum systems we now consider are the following: the first one is a generalised oscillator 444 Here we consider two different cases, namely when a constant driving force is present and the other one is when there is no driving force. and the second one is an isotonic oscillator where, apart from a simple harmonic term, an additional singular inverse square term is also present in the total potential. The CG generated by such Hamiltonians can be obtained by using the zeta-function regularization technique and their implications will be discussed. All the systems considered in this paper are one-dimensional.
@@ III.1 Complexity geometry generated by a generalised oscillator Hamiltonian
III.2 Complexity geometry generated by the generalised oscillator Hamiltonian with a constant driving force
which, as the subscript above suggests, is just that of a HO with a constant term added. Notice that this constant term actually shifts the energy eigenvalues of the system, so that the analogues of the relations in Eq. (12) and Eq. (22) can be written without introducing the constant $a$ by hand, and are given by,
III.3 Complexity geometry generated by the isotonic oscillator Hamiltonian
The norm of the Hamiltonian, in turn, is given by
III.4 Expression for the complexity
@@ IV Complexity geometry of a harmonic oscillator with time-dependent frequency
From this section onward, we shall consider Hamiltonians, which are explicitly time-dependent. Our essential goal is to calculate the regularised cost function of the form $F^{2}_{r}\big{(}\mathbf{H}(t),t\big{)}=\lambda_{0}^{2}\text{Tr}\big{[}\big{(}\mathbf{H}(t)-a(t)\mathbf{I}\big{)}^{2}\big{]}$. Here we have introduced an additional argument $t$ in the expression for the cost to explicitly indicate that it is a function of time characterising the history of the evolution. We first obtain the diagonal and off-diagonal matrix elements of a time-dependent Hamiltonian in a suitable orthogonal basis, and use this to obtain the norm of the Hamiltonian which is the bi-invariant cost in subsection IV.1 and IV.2 respectively. This involves a regularisation procedure similar to the one discussed in the previous section, which we elaborate on in subsection IV.3. We finish this section with two illustrative analytical examples in IV.4 and IV.5 where the exact expression for the regularised cost can be written down in compact form.
@@ IV.1 Matrix elements of the Hamiltonian
where we have taken the mass of the oscillator to be unity. Our goal is to obtain the bi-invariant CG generated by the Hamiltonian of the OTF, i.e., we need to calculate the trace of the displaced time-dependent Hamiltonian. As before, we can add a pure time-dependent term with the above Hamiltonian from the beginning, and interpret it as the mean value used to regularise the trace of this Hamiltonian, and since it only adds a time-dependent pure phase factor in the total wavefunction, the displaced Hamiltonian $\tilde{\mathbf{H}}_{\Omega}(t)$ and the original one $\mathbf{H}_{\Omega}(t)$  [^5] From now on, in the notation for the Hamiltonian, we shall only show its explicit dependence on time and suppress other operators for convenience. correspond to the same CG. The classical equation of motion derived from the above Hamiltonian is given by the usual second order equation

### v2

@@ I Introduction
The quantity computational complexity of a given quantum state roughly measures the minimum number of basic gates required to build the target state starting from an initial reference state. Among various measures to quantify computational complexity, one of the most commonly used is the one proposed by Nielsen and collaborators Nielsen1 ; Nielsen2 ; Nielsen3 . In this approach, the reference and target states correspond to two points on a curved manifold, and all the possible unitary transformations that connect these two points are assigned with a cost function. According to this geometric formulation, the path that minimises the cost function is nothing but the geodesic on that manifold of the unitary operator, and the length of the geodesic is called the Nielsen complexity (NC). This particular measure of complexity has been explored in significant detail in recent literature not only because it can shed light on fundamental aspects of the information contained in a quantum state Myers1 ; BSS ; KKS ; Myers2 ; HM ; BMS ; BS ; Caceres:2019pgf , but also due to the fact that NC can also be used an important probe of quantum quenches ABHKM ; Gautam:2022gci ; Pal:2022rqq ; Chandran:2022vrw ; Camargo:2018eof , quantum phase transitions LWCLTYGG ; JGS ; XYZ ; SK ; JGS2 ; PPS ; WH ; Roca-Jerat:2023mjs ; Jaiswal:2024yts ; Jayarama:2022xta , and quantum chaos BDKP1 ; BDKP2 ; HJU ; Qu:2022zwq , and has various other applications as well Pal:2022ptv ; Chowdhury:2023iwg ; Adhikari:2021ckk ; Basteiro:2021ene ; Camilo:2020gdf ; Erdmenger:2020sup . Specifically, typical behaviour of circuit complexity has been used in ABHKM ; Camargo:2018eof ; Pal:2022rqq ; Gautam:2022gci ; Chandran:2022vrw to understand the unitary dynamics of an isolated quantum system, when the system is taken out of equilibrium by means of a single or multiple sudden or smooth quenches thereby changing the parameters of the Hamiltonian according to different relevant time scales. It has also been used in various works as an important quantum information theoretic tool to detect the zero temperature quantum phase transition in various contexts and models, where typically the discontinuity of the complexity marks the quantum critical point LWCLTYGG ; JGS ; XYZ ; SK ; JGS2 ; PPS ; WH ; Roca-Jerat:2023mjs ; Jaiswal:2024yts ; Jayarama:2022xta . Apart from the NC, there are also other important measures of circuit complexity of quantum states which differ from Nielsen’s approach in many aspects (e.g., the one introduced in CHMP by using the Fubini-Study metric associated with a pure quantum state), however, in this paper, we shall mostly consider Nielsen’s formulation. For a recent review of different approaches to circuit complexity and its application in holography, we refer to Chapman:2021jbh .
@@
It is important to note that, apart from the given target and reference states, the NC crucially depends on the choice of the cost function associated with each of the unitary paths on the manifold. Though the qualitative nature of NC in a quantum system strongly depends on the choice cost function, in Nielsen’s original formulation, this choice remains arbitrary (for an exposition on different cost functions used in this context, we refer to BMS ). Another built-in aspect of NC is the fact that the unitary complexity geometry associated with NC is either right or left-invariant under a global transformation of the generator of the unitary curve connecting initial and final operators. As was pointed out in YK ; YK2 ; YANZK ; YANZK2 , for a physical quantum system, when this generator is the Hamiltonian of the system itself, the lack of unitary invariance is somewhat unsatisfactory since, e.g., the Nielsen complexity geometry corresponding to two Hamiltonians related by a unitary transformation would be different, and hence unacceptable from a physical point of view since they describe the same quantum system. Based on the assumption that NC and hence the cost function should be invariant under both left and right unitary transformation of the generator associated with a given path on the unitary manifold, the notion of bi-invariant complexity was proposed in YK ; YANZK ; YANZK2 . In this paper, we mainly consider this bi-invariant formulation of the NC, and apply it to the Hamiltonians of some well-known quantum mechanical systems.
@@
Usually, in the problem of calculating NC (or other notions of circuit complexity), the following situation is considered: the reference and target quantum states are specified (along with the chosen cost functional and penalty factors for any gate directions, if any), and one has to minimise the costs associated with all the relevant unitary curves to find the NC. However, in this approach the instantaneous generator of the unitary curve is not usually given specific importance since one is only interested in the length of the geodesic curve (i.e., complexity) between two states (or operators) on the unitary complexity geometry (CG), irrespective of the generator of that path. Consider now a different posing of the above-mentioned problem, which is equally important: the time evolution of an initial reference under the Hamiltonian of a given quantum system. In that case, the unitary operator is the time evolution operator itself, and the target state is actually some time-evolved state for a particular value of time. Of course, any arbitrary target state may not be on the unitary curve generated by the system Hamiltonian, and consequently, those states are not achievable under this protocol.
From a different perspective, suppose that one can choose only two from the following three quantities: a reference state, a target state, and the (Hermitian) generator of the path connecting these two states on a CG, while the third one is naturally constrained by the unitary evolution. In the standard way of defining NC, the first two (i.e., reference and target states) are given, whereas here we are interested in the scenario where the first and third options (i.e., reference state and system Hamiltonian) are pre-assigned. As a natural consequence of the preceding statement, the nature of the parameter used to parametrise the unitary curve connecting the reference and target state is different from the usual treatment of NC. Specifically, in this paper, we construct a bi-invariant CG associated with the operators generated by the Hamiltonian time evolution of quantum mechanical systems. Therefore, in that case, the parameter along the unitary path connecting the reference and target state is time. And the reason for choosing the bi-invariant geometry for this purpose is the fact that, on the one hand, it is invariant under unitary transformations which leave the Hamiltonian invariant, while on the other hand, since this geometry solely depends on the generator of the unitary evolution curve, i.e., here the Hamiltonian of the quantum system under consideration, its computation is relatively straightforward for simple systems. In summary, our goal here is to calculate the cost of operators along the time evolution curve of a quantum mechanical system by using the framework of bi-invariant CG.
@@
In Section V, we generalise the results of section IV for the case of the Hamiltonian of an oscillator whose both mass and frequency changes with time. For this Hamiltonian also, we are able to provide an analytical formula for the regularised cost function by using the Lewis - Riesenfeld theory of time-dependent invariant operators. The fact that such a Hamiltonian is actually an element of the $su(1,1)$ Lie algebra provides a way of directly obtaining the exact form of the time evolution operator using techniques of Lie group theory. Using this exact expression for the time-evolution operator, we show that suitably redefining the time parameter, it is possible to map the time evolution operator of an oscillator with time varying mass and frequency to an oscillator with unit mass but time varying frequency. Furthermore, we show that under the same reparametrisation, the total bi-invariant cost evaluated along the time-evolution curves of these two oscillators can be mapped to each other. In section VI we compute the bi-invariant cost of a oscillator with time-dependent frequency and compare it with other measures, such as the change in Shannon entropy in one of the most common protocols where such time-dependent Hamiltonians arise, namely that of a smooth quantum quench.
We discuss the significance of our results in section VII, and indicate a few possible future directions. This paper also contains three appendices; in Appendix. A and B, we discuss further examples of calculation of the bi-invariant geometry of, respectively, time-independent and explicit time-dependent Hamiltonians, whereas in Appendix. C we provide some details of the computation of the time-evolution operator associated with Hamiltonians which are elements of the $su(1,1)$ Lie algebra.
@@ II Cost functional and the circuit complexity in Nielsen’s approach
Now we go back to the problem of defining such a geometry uniquely. Besides the above mentioned problem of nonuniqueness of the CG, there is one additional feature of such constructions - the resulting geometry is only right (or left) invariant under a global transformation of the unitary operator. As discussed in detail in YANZK ; YANZK2 ; YK , this property is somewhat problematic for defining complexity associated with a quantum mechanical Hamiltonian, since a unitary transformed system is identical to that of the original system, in the sense that physical observables computed from both of them make identical predictions for the original system and the transformed one. Hence, it is very much desirable that for an arbitrary unitary operator $U(s)$, defined on the CG, the cost function or the norm should be such that it is invariant under the unitary transformation of the Hamiltonian, i.e.,
@@ III Complexity geometry associated with time-independent Hamiltonians
We now discuss the construction of the bi-invariant CG of operators generated by the Hamiltonian of some well-known quantum systems. In this section we consider the cases where the Hamiltonian is time-independent, while in the next section, and the rest of the paper, we deal with systems whose Hamiltonians are explicit functions of time. The instantaneous Hamiltonian appearing in Eq. (2) is the ordered Hermitian Hamiltonian operator (which we denote as $\mathbf{H}$ in the followings) of a given quantum system, and this equation is nothing but the Schrodinger equation for the time evolution of the state $\big{|}\psi(t)\big{>}$. Furthermore, the unitary operator of our concern is the time evolution operator generated by the system Hamiltonian. At this point, we stress that we want to evaluate the computational cost along the time evolution curve $\mathcal{U}(t)$, not the NC of a time evolved state (or an operator on the time evolution curve) - this fact makes our analysis different from other works where time evolution of NC corresponding to different quantum systems have been reported BDKP1 ; YK2 ; HJU . Specifically, in our set-up, the parameter $s$ characterising the curves in the space of unitary operators is time $t$.  [^1] In this sense, the motivation of the present paper is similar to the recent set of works Craps:2022ese ; Craps:2023ivc ; Craps:2023rur , which has obtained an upper bound on the NC by restricting the minimisation procedure described in the previous section within a prescribed family of curves connecting the identity and the target operator.
@@
In that case, the bi-invariant cost represents a length on a Riemannian metric, which is nothing but the Cartan-Killing metric corresponding to the Lie group whose generator is the Hamiltonian of the quantum system under consideration RG ; FSS . For the systems we consider in this paper, the Hamiltonian is an element of some Lie algebra, and the above mentioned Lie group is associated with this algebra. Clearly, to find out the CG using this cost, we need to calculate the trace of the squared Hamiltonian. However, the Hilbert spaces associated with the systems we are going to consider are infinite dimensional, and hence the trace can be divergent. A procedure of obtaining the CG by suitably regularizing the trace of the Hamiltonian operator using the Hurwitz $\zeta$ functions was described in YK . For later purposes, we first briefly review this procedure and its application to find out the CG of a generalised oscillator, as worked out in YK . We can also provide an alternative interpretation of this regularization in terms of a constant driving force present in the original Hamiltonian (see subsection A.1).
@@
To illustrate the above procedure, the quantum systems we consider are the following: the first one is a generalised oscillator 444 We consider two different cases, namely when a constant driving force is present (which is discussed in Appendix. A in detail) and the other one is when there is no driving force. The bi-invariant cost function associated with a generalised oscillator without a constant driving force was obtained YK . Here we briefly review this procedure which we shall then generalise in the next section for time-dependent Hamiltonians. and the second one is an isotonic oscillator where, apart from a simple harmonic term, an additional singular inverse square term is also present in the total potential (this example is discussed in Appendix A). The CG generated by such Hamiltonians can be obtained by using the zeta-function regularization technique and their implications will be discussed. All the systems considered in this paper are one-dimensional.
@@ III.1 Complexity geometry generated by a generalised oscillator Hamiltonian
III.2 Expression for the complexity
@@ IV Complexity geometry of a harmonic oscillator with time-dependent frequency
From this section onward, we shall consider Hamiltonians which are explicitly time-dependent. This kind of Hamiltonian commonly appears when the quantum system under consideration is influenced by some external perturbation (i.e., are not closed) so that the parameters in the Hamiltonian become functions of time. One example of such a system can be an atom undergoing radiative transitions such that its interaction with the classical radiation field contributes an explicitly time-dependent term in the Hamiltonian. For the purposes of the present paper, the situations that are relevant are the systems that are driven far from equilibrium by some external perturbation, a process commonly known as the quantum quenches. When this process occurs during a finite interval of time, the Hamiltonian governing the time evolution of the system becomes explicitly time-dependent. See section VI for an example of such quench protocol where we shall explicitly compute the cost derived later in this section and compare it with other information theoretic quantities, such as change in the Shannon entropy.
Our essential goal in this section is to calculate the regularised cost function of the form $F^{2}_{r}\big{(}\mathbf{H}(t),t\big{)}=\lambda_{0}^{2}\text{Tr}\big{[}\big{(}\mathbf{H}(t)-a(t)\mathbf{I}\big{)}^{2}\big{]}$. Here we have introduced an additional argument $t$ in the expression for the cost to explicitly indicate that it is a function of time characterising the history of the evolution. We first obtain the diagonal and off-diagonal matrix elements of a time-dependent Hamiltonian in a suitable orthogonal basis, and use this to obtain the norm of the Hamiltonian which is the bi-invariant cost in subsection IV.1 and IV.2 respectively. This involves a regularisation procedure similar to the one discussed in the previous section, which we elaborate on in subsection IV.3. We finish this section with two illustrative analytical examples in IV.4 and IV.5 where the exact expression for the regularised cost can be written down in compact form.
@@ IV.1 Matrix elements of the Hamiltonian
where we have taken the mass of the oscillator to be unity. We also assume that the time-dependent frequency $\Omega(t)$ is always a real function of time. Our goal is to obtain the bi-invariant CG generated by the Hamiltonian of the OTF, i.e., we want to calculate the trace of the displaced time-dependent Hamiltonian. As before, we can add a pure time-dependent term with the above Hamiltonian from the beginning, and interpret it as the mean value used to regularise the trace of this Hamiltonian, and since it only adds a time-dependent pure phase factor in the total wavefunction, the displaced Hamiltonian $\tilde{\mathbf{H}}_{\Omega}(t)$ and the original one $\mathbf{H}_{\Omega}(t)$  [^5] From now on, in the notation for the Hamiltonian, we shall only show its explicit dependence on time and suppress other operators for convenience. correspond to the same CG. The classical equation of motion derived from the above Hamiltonian is given by the usual second order equation